# Meta-DAN: towards an efficient prediction strategy for page-level handwritten text recognition

### v1

@@ V-B Training strategy
  • We pre-train the encoder part of the architecture using a line-level training strategy. An adaptive max-pooling layer is added on top of the encoder to collapse the vertical axis, leading to a one-dimensional sequence of features. Then classification layer is applied to generate a probability lattice. This pre-training is carried out using the CTC loss. If line-level annotations are provided in the dataset, we use a ratio of 25% real / 75% synthetic images (100% synthetic otherwise).
@@
  • We randomly apply some data augmentation strategies with a probability of 90%. It includes resolution modification, color jittering, Gaussian noising, Gaussian blurring, elastic distortion, dilation, erosion and perspective transformation.
@@
  • During page-level training, we rely on a curriculum synthetic document generation strategy. The generated documents are more and more complex through the training, starting with single line images to documents including several dozens lines. The model is trained on both real and synthetic documents with the proportion of synthetic documents slowly decreasing from 90% to 10%. Following [17], we scraped the website 1001-fonts.com to get a large collection of fonts from which we randomly sample when rendering a synthetic text line. The text used for synthetic generation comes from the training set of the target dataset itself.
@@
  • Training and pre-training are performed using the Adam optimizer with an initial learning rate of $10^{-4}$, and automatic mixed-precision.
  • Line-level pre-training is carried out using 5M training samples, and page-level training (including curriculum) is performed over 1M samples.
  • Input images are preprocessed to reach a resolution of 150 dpi by preserving the original aspect ratio, whether it is for pre-training, training or evaluation.
  • We do not use any external data, external language model nor lexicon constraints.
V-C Evaluation
@@
We use the standard Character Error Rate (CER) and the Word Error Rate (WER) to evaluate the text recognition performance. Both are computed with the Levenshtein distance between the ground truth and the predicted text, normalized by the length of the ground truth. It has to be noted that page-level approaches are sensitive to reading order; it means that correct text recognized in a wrong order will be considered as incorrect. For each model, we compute the average prediction time for one test image when using a single A100 GPU and a mini-batch size of 1. To provide fully comparable experiments, notably between monolingual and multilingual experiments we present in the following, we set the same maximum number of decoding iterations for all datasets. It is set to 5,000 to cover the requirements of all evaluated dataset. It has to be noted that this choice results in recognition metrics that can highly be impacted by the prediction loops that can occur, i.e., when the model predicts again and again the same text line.
VI Experiments
@@
In this section, we evaluate the proposed approaches (W-DAN, MT-DAN and Meta-DAN) on the IAM and BRESSAY datasets to select the best approach and hyperparameters. While relying on the original DAN architecture [13], we propose an enlarged version of the encoder we refer to as FCN-M as it shows better recognition performance (details in Appendix A). Then, we apply them on the other 9 datasets to validate the robustness of the approach, and compare it to the state of the art. Finally, we present the results for a multilingual recognition setting.
VI-A W-DAN
@@
We evaluate the W-DAN approach for window sizes $w\in\{1,2,3,4,5,10,25\}$, with $w=1$ corresponds to the DAN setting. Results are shown in Table V-A. It highlights that it is possible to predict from 2 to 10 character tokens at once, while still being competitive in terms of CER and WER. It is especially true for the BRESSAY dataset for which $w=10$ leads to a small increase of only 0.23 point of CER, for an 8.4 speed factor. However, results degrade significantly for $w=25$ with a CER of 6.56% instead of 4.23% for IAM, and 4.37% instead of 2.43% for BRESSAY, even with three times more training samples. As one can note, it is possible to reduce the average prediction time per sample nearly linearly with respect to the window size. This is illustrated in Figure VI-A. Regarding the BRESSAY prediction time, it appears that the new implementation has impacted the prediction time for large text predictions, leading to a lag in the speed factor improvement. A main observation is that the model is harder and harder to train as $w$ increases, so we needed to increase the amount of training data for $w=10$ and $w=25$ up to 2M and 3M, respectively.
Evolution of the prediction time with respect to the window size $w$ for the W-DAN and to the number of token heads $m$ for the MT-DAN on the test set of IAM and BRESSAY datasets. The prediction time dramatically decreases as $m$ or $w$ increases at first, and then less sharply as the computation load of the encoder becomes more and more significant on the overall inference process.
VI-B MT-DAN
@@
[b] Comparison of MT-DAN recognition performance with respect to the number of token heads $m$ on the IAM and BRESSAY test sets. Speed Factor (SF) is indicated as a comparison with the DAN baseline. Number of IAM BRESSAY heads ($m$) CER (%) WER (%) Time (s) SF CER (%) WER (%) Time (s) SF 1 (DAN) 4.23 14.00 3.39 x1.0 2.43 7.85 24.65 x1.0 2 4.70 14.85 1.84 x1.8 2.47 8.54 13.02 x1.9 3 4.74 15.17 1.23 x2.8 2.46 8.60 8.85 x2.8 4 4.68 15.56 0.93 x3.6 2.70 8.91 6.64 x3.7 5 5.14 16.01 0.75 x4.5 2.45 8.72 5.28 x4.7 10* 5.06 16.29 0.39 x8.7 2.37 8.64 2.64 x9.3 25* 7.19 23.83 0.16 x21.2 3.15 11.28 1.11 x22.2
    Training was performed over 2M training samples for $m=10$ and $m=25$. An additional character-level curriculum stage was also added for the early stages of the training.
@@
We evaluate the MT-DAN approach in a similar setting, i.e., for various numbers of token heads $m\in\{1,2,3,4,5,10,25\}$, with $m=1$ corresponds to the original DAN. The obtained results (Table VI-B) are quite similar to those of the W-DAN, with comparable recognition performance and prediction time. Neither approach seems to stand out from the other. Even for the extreme case of $m=25$, it appears that the MT-DAN performs better for the BRESSAY dataset with a CER of 3.15% compared to 4.37% for the W-DAN (with $w=25$), but we observe the contrary for the IAM dataset with a CER of 7.19% for the MT-DAN compared to 6.56% for the W-DAN. As for the W-DAN, it required more training data for $m=10$ and $m=25$. We also relied on a curriculum strategy at character level before the line-level curriculum strategy takes place. We did not use it for the W-DAN as it was not effective; we assume this is due to the way the tokens interact for make its predictions in this approach.
@@
To enhance the capacity of the MT-DAN approach, we propose and evaluate two prediction policies at inference time:
@@
1) The static prediction strategy consists in always predicting the $k$ next tokens at each decoding iteration ($k$ must be lower or equal to the number of tokens heads $m$). Previous results (shown in Table VI-B) relied on this naïve prediction policy, setting $k=m$.
@@
2) The dynamic prediction strategy involves predicting between the next 1 and $m$ tokens, based on a confidence threshold $\tau$. This way, the number of predicted tokens must vary from one decoding iteration to another. The idea is to accept to slow down the decoding process when the model is unsure of its prediction, by assuming that it could be better at the next decoding iteration for which it will benefit from more context. The first token head is always used to avoid infinite loop; the other $m-1$ heads are then considered sequentially. If the confidence score (i.e., the maximum value after softmax operation) associated to the current head is greater or equal to $\tau$, then the prediction is preserved, and the next head is processed; the decoding iteration ends otherwise.
We first study the choice of the confidence threshold $\tau$ for the dynamic strategy. Figure VI shows the evolution of the CER with respect to $\tau$ for values between 0% and 100%, with a step of 5%. We compare it for different numbers of heads and for both IAM and BRESSAY datasets. It appears that the recognition performance can globally be improved as of $\tau=0.5$, except for the configurations $m=3$ and $m=4$ on the BRESSAY dataset whose behavior is less monotonous. One should note that this prediction policy can even reach better CER performance than the DAN (blue line) on the BRESSAY dataset, for $w=2$ and $w=5$. We consider $\tau=0.9$ as a good compromise over all the evaluated configurations and use it for the next experiments.
Comparison of static and dynamic prediction policies of the MT-DAN. CER (%) are given for the test sets of IAM and BRESSAY datasets. Best results are shown in bold; results that are very close to the best (whose difference in values is less than 0.05%) are underlined. Number IAM BRESSAY of heads static dynamic static dynamic ($m$) $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $\tau=0.9$ time (s) $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $\tau=0.9$ time (s) 1 (DAN) 4.23 2.43 2 4.33 4.70 4.59 1.92 2.41 2.47 2.36 13.60 3 4.31 4.85 4.74 4.31 1.32 2.46 2.46 2.46 2.69 9.21 4 4.48 4.54 4.68 4.68 4.52 1.00 2.71 2.86 2.61 2.70 2.71 6.97 5 4.45 4.74 4.54 4.86 5.14 4.51 0.83 2.34 2.40 2.36 2.40 2.45 2.33 5.70 10 4.79 4.80 4.87 4.87 4.90 4.80 0.47 2.67 2.24 2.70 2.28 2.29 2.66 3.15 25 5.73 5.61 5.58 5.64 5.68 5.47 0.36 3.35 3.30 2.92 2.91 2.93 2.87 1.88
Impact of the CER threshold for dynamic decoding of the MT-DAN on the test set of IAM and BRESSAY.
@@
We compare the two prediction policies, namely static and dynamic, in Table VI. For the static prediction strategy, we evaluate the models for $k\in\{1,2,3,4,5\}$ when possible. Regarding the IAM dataset, one can observe a trend whereby the character error rate decreases as $k$ also decreases, with optimum performance reached for $k=1$, noting that decreasing $k$ highly increases the prediction time as shown previously. This trend is less noticeable on the BRESSAY dataset with more fluctuating results. However, the proposed dynamic prediction policy with $\tau=0.9$ shows interesting results which generally outperforms the corresponding model using the static strategy with $k=m$, with a limited prediction time overhead. For instance, it enables to reduce the CER from 5.14% to 4.51% for the IAM dataset ($w=5$), while slightly increasing the prediction time from 0.75 second (Table VI-B) to 0.83 second.
VI-C Meta-DAN
In this section, we focus on $m=5$ and $w=5$ as it seems to be a good comprise between prediction time and recognition performance.
Pre-training strategy
Comparison of Meta-DAN ($m=5$, $w=5$) pre-training strategy on the test sets of IAM and BRESSAY. Pre-training curriculum IAM BRESSAY line-level (baseline) 5.59 3.01 W-DAN ($w=5$) 4.82 2.51 W-DAN ($w=5$) 5.01 2.42 MT-DAN ($m=5$) 5.04 2.36 MT-DAN ($m=5$) 4.79 2.38
In Table VI, we evaluate the Meta-DAN for several pre-training strategies. We compare the standard line-level pre-training to transfer learning directly from page-level W-DAN or MT-DAN, with and without curriculum learning. We observe that best results are obtained through page-level pre-training, highlighting the usefulness of the pre-trained decoder even if it does not rely on the same decoding strategy. We also note that re-training from line-level through curriculum learning is generally not necessary to reach the best results. Lastly, it seems that the MT-DAN pre-training leads to slightly better CER on average. For these reasons, the Meta-DAN training will be based on MT-DAN pre-training without curriculum learning in the following experiments.
@@ VI-C Meta-DAN
We compare the Meta-DAN with the MT-DAN and the W-DAN for the same hyperparameters ($m=5$ and $w=5$) in Table VI. As one can note for the IAM and BRESSAY datasets, combining both windowed queries and multi-token predictions enable to increase the recognition performance, as well as reducing the prediction time. Comparing to DAN, the conclusion is not as clear-cut regarding the character error rate as it is in favor to the DAN for IAM with a CER of 4.23% compared to 4.79%, but in favor of the Meta-DAN for BRESSAY with a CER of 2.38% compared to 2.43%. It shows that the Meta-DAN can be both more efficient and faster than the DAN depending on the target dataset.
@@
Comparison of proposed approaches using the static prediction policy (with $m=k$). CER are expressed in percentages, and times in seconds. Approach $m$ $w$ IAM BRESSAY CER Time CER Time DAN 1 1 4.23 3.39 2.43 24.65 W-DAN 1 5 4.87 0.77 3.02 6.63 MT-DAN 5 1 5.14 0.75 2.45 5.28 Meta-DAN 5 5 4.79 0.43 2.38 4.01
Comparing prediction strategies
As the Meta-DAN also relies on multi-token predictions, we can apply the same dynamic prediction policy as for the MT-DAN. Its usage is evaluated in Table VI. It appears that the confidence scores associated to the predictions are not as high as for the MT-DAN, leading to a failure case for the BRESSAY dataset for which the prediction time is close to that of W-DAN (static case with $k=1$). As noted previously, the static policy with $k=1$ leads to the best recognition performance but does not benefit from accelerated prediction. We consider the static policy with $k=m=5$ as the best alternative for the Meta-DAN and use this configuration in the following.
@@
Comparison of Meta-DAN ($w=5$, $m=5$) performance with respect to the decoding strategy on the test sets of IAM and BRESSAY. Strategy for IAM BRESSAY concurrent CER (%) Time (s) CER (%) Time (s) static ($k=1$) 4.55 0.79 2.31 7.16 static ($k=5$) 4.79 0.43 2.38 4.01 dynamic ($\tau=0.9$) 4.73 0.45 2.37 7.14
VI-D Visualization

### v2

@@ 5.2 Training strategy
  • The encoder is pre-trained using a CTC-based line-level training strategy. An adaptive max-pooling layer is added on top of the encoder to collapse the vertical axis, leading to a one-dimensional sequence of features. Then, a classification layer is applied to generate a probability lattice. If line-level annotations are provided in the dataset, we use a ratio of 25% real / 75% synthetic images (100% synthetic otherwise).
@@
  • Data augmentation is applied with a probability of 90%, including resolution change, color jittering, Gaussian noising, Gaussian blurring, elastic distortion, dilation, erosion, and perspective transformation.
@@
  • The page-level training relies on a curriculum synthetic document generation strategy. The generated images contain more and more text lines through training. The ratio of synthetic documents slowly decreases from 90% to 10%. Following [11], we rely on the 1001-fonts.com website to get a large collection of fonts from which we randomly sample when rendering a synthetic text line. The text used for synthetic generation is from the training set of the target dataset.
@@
  • Line-level pre-training is carried out using 5M training samples; page-level training (including curriculum) is performed over 1M samples. Training and pre-training are performed using the Adam optimizer with an initial learning rate of $10^{-4}$, and automatic mixed-precision.
  • Preprocessing: images are scaled to reach 150 dpi resolution. No post-processing: we do not use any external data or language model.
5.3 Evaluation
@@
We compute the standard Character Error Rate (CER) and Word Error Rate (WER) to evaluate the text recognition performance. Both metrics are computed using the Levenshtein distance [29], which counts the number of deletion, addition, and substitution operations required to go from the prediction sequence to the ground truth sequence, normalized by the length of the ground truth sequence. While the CER considers the prediction and the ground truth as sequences of characters, the WER considers them as sequences of words. All results are given for the test sets. It has to be noted that page-level approaches are sensitive to reading order; it means that correct text recognized in a wrong order will be considered as incorrect. For each model, we compute the average prediction time for one test image when using a single A100 GPU and a mini-batch size of 1. We set the maximum number of decoding iterations for all datasets to 5,000 to cover the requirements of all evaluated datasets. This way, we do not rely on dataset-specific properties, which is closer to a real-case scenario. It has to be noted that this choice results in recognition metrics that can be highly impacted by the prediction loops that can occur, i.e., when the model predicts again and again the same text line.
6 Experiments
@@
In this section, we evaluate the proposed approaches (W-DAN, MT-DAN, and Meta-DAN) on the IAM and BRESSAY datasets to select the best approach and hyperparameters. Then, we apply them to the other 9 datasets to validate the robustness of the approach. While relying on the original DAN architecture [14], we propose an enlarged version of the encoder, which we refer to as FCN-M as it shows better recognition performance (details in B).
6.1 W-DAN
Table 6.1(a) provides the results of the W-DAN approach for several window sizes $w$. It highlights that it is possible to predict up to 10 character tokens at once, while being competitive in terms of CER. It is especially true for the BRESSAY dataset for which $w=10$ leads to a small increase of only 0.23 points of CER. However, results degrade significantly for $w=25$ with a CER of 6.56% instead of 4.23% for IAM, and 4.37% instead of 2.43% for BRESSAY.
@@ 6.1 W-DAN
As one can note, it is possible to reduce the average prediction time per sample nearly linearly with respect to the window size. Regarding BRESSAY prediction times, it appears that the new implementation has impacted the prediction time for large text predictions, leading to a lag in the speed factor improvement. A main observation is that the model is harder to train as $w$ increases. We assume that this is due to the inherent complexity of the self-attention layer, which grows with the length of the query sequence. Indeed, the more query tokens are processed altogether, the greater the required effort to coordinate and focus on the right parts of the image features. In this way, we needed to increase the amount of training data to reach convergence for $w=10$ and $w=25$ up to 2M and 3M, respectively.
@@
Comparison of W-DAN and MT-DAN performance with respect to the window size $w$ and to the number of token heads $m$, respectively, on the IAM and BRESSAY test sets.
[b]
(a) W-DAN results.
$w$ IAM BRESSAY
CER (%) Time (s) CER (%) Time (s)
1 (DAN) 4.23 3.39 2.43 24.65
2 4.67 1.94 2.91 20.86
3 4.83 1.26 3.05 12.25
4 4.75 0.97 2.63 8.63
5 4.87 0.77 3.02 6.63
10* 5.07 0.40 2.66 2.95
25* 6.56 0.18 4.37 1.13
(b) MT-DAN results.
$m$ IAM BRESSAY
CER (%) Time (s) CER (%) Time (s)
1 (DAN) 4.23 3.39 2.43 24.65
2 4.70 1.84 2.47 13.02
3 4.74 1.23 2.46 8.85
4 4.68 0.93 2.70 6.64
5 5.14 0.75 2.45 5.28
10* 5.06 0.39 2.37 2.64
25* 7.19 0.16 3.15 1.11
    More training data was used for $w=10$, $m=10$, $w=25$, and $m=25$
6.2 MT-DAN
@@
For fair comparison, we evaluate the MT-DAN approach for the same values of token heads $m\in\{1,2,3,4,5,10,25\}$, as both $m$ for MT-DAN and $w$ for W-DAN refer to the number of predicted tokens per decoding iteration. Here, $m=1$ corresponds to the original DAN. The obtained results shown in Table 6.1(b) are quite similar to those of the W-DAN, with comparable recognition performance and prediction time. Neither approach seems to stand out from the other. Even for the extreme case of $m=25$, it appears that the MT-DAN performs better for the BRESSAY dataset with a CER of 3.15% compared to 4.37% for the W-DAN (with $w=25$), but it is the contrary for the IAM dataset: 7.19% of CER compared to 6.56% for the W-DAN. As for the W-DAN, it required more training data for $m=10$ and $m=25$ to reach convergence. However, this is not for the same reason. We assume that, for the MT-DAN, this is due to the compression task that is harder as $m$ increases. Indeed, $m$ character representations must be compressed into a single fixed-size vector. We also relied on a curriculum strategy at character level before the line-level curriculum strategy. We did not use it for the W-DAN as it was not effective; we assume this is due to the way the tokens interact to make their predictions in this approach.
@@
To enhance the capacity of the MT-DAN approach, we propose and evaluate two variants, we called prediction policies, at inference time:
@@
1) The static prediction policy consists in always predicting the $k$ next tokens at each decoding iteration ($k\leq m$). Previous results, shown in Table 6.1(b), relied on this naïve prediction policy, setting $k=m$.
@@
2) The dynamic prediction policy involves predicting between $1$ and $m$ tokens at once, based on a confidence threshold $\tau$. The number of predicted tokens must vary from one decoding iteration to another. The idea is to predict fewer tokens when the model is unsure of its prediction, by assuming that it could be better at the next decoding iteration, for which it will benefit from more context. The first token head is always used to avoid an infinite loop; the other m1m-1 heads are then considered sequentially. If the confidence score (i.e., the maximum value after softmax operation) associated with the current head is greater than or equal to $\tau$, then the prediction is preserved, and the next head is processed; the decoding iteration ends otherwise.
Accuracy-speed trade-off with respect to confidence threshold $\tau$ for dynamic decoding of the MT-DAN on IAM and BRESSAY.
We first study the choice of the confidence threshold $\tau$ for the dynamic strategy. Figure 6.2 shows the evolution of the CER and of the prediction time with respect to $\tau$ for values between 0 and 1, with a step of 0.1. A confidence threshold of $\tau=0$ means that the token predicted by each projection head is always kept (the fastest option): it corresponds to the static prediction policy. A confidence threshold of $\tau=1$ means that only the prediction of the first projection head is kept (the longest option): it results in the DAN prediction strategy. We compare the accuracy-speed trade-off for different numbers of heads on both IAM and BRESSAY datasets. We also show results for a single projection head as a baseline, leading to a CER plateau as the dynamic strategy is designed for multi-token predictions.
It appears that the confidence threshold fulfills its purpose, as there is a trend toward a decrease in the CER as $\tau$ increases, for both datasets. In addition, the prediction time per sample is rather stable from $\tau=0$ to $\tau=0.9$. However, one can note a huge prediction time gap going from $\tau=0.9$ up to $\tau=1$, making the approach ineffective as it consists in single-token prediction, as the DAN. In this way, we consider $\tau=0.9$ as a good compromise over all the evaluated configurations and use it for the next experiments.
@@
We compare the two prediction policies, namely static and dynamic, in Table 6.2. For the static prediction policy, we evaluate the models for $k\in\{1,2,3,4,5\}$ when possible. Regarding the IAM dataset, one can observe a trend whereby the CER decreases as $k$ also decreases, with optimum performance reached for $k=1$, noting that decreasing $k$ highly increases the prediction time. This trend is less noticeable on the BRESSAY dataset with more fluctuating results. However, the proposed dynamic prediction policy with $\tau=0.9$ shows interesting results, which generally outperform the corresponding model using the static strategy with $k=m$, with a limited prediction time overhead. For instance, it enables to reduce the CER from 5.14% to 4.51% for the IAM dataset ($m=5$), while slightly increasing the prediction time from 0.75 seconds (Table 6.1(b)) to 0.83 seconds.
Comparison of static and dynamic prediction policies of the MT-DAN. CER (%) are given for IAM and BRESSAY datasets. Best results are shown in bold; results that are very close to the best (whose difference in values is less than 0.05%) are underlined. Number IAM BRESSAY of heads static dynamic static dynamic ($m$) $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $\tau=0.9$ Time $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $\tau=0.9$ Time 1 (DAN) 4.23 2.43 2 4.33 4.70 4.59 1.92 2.41 2.47 2.36 13.60 3 4.31 4.85 4.74 4.31 1.32 2.46 2.46 2.46 2.69 9.21 4 4.48 4.54 4.68 4.68 4.52 1.00 2.71 2.86 2.61 2.70 2.71 6.97 5 4.45 4.74 4.54 4.86 5.14 4.51 0.83 2.34 2.40 2.36 2.40 2.45 2.33 5.70 10 4.79 4.80 4.87 4.87 4.90 4.80 0.47 2.67 2.24 2.70 2.28 2.29 2.66 3.15 25 5.73 5.61 5.58 5.64 5.68 5.47 0.36 3.35 3.30 2.92 2.91 2.93 2.87 1.88
6.3 Meta-DAN
In this section, we focus on $m=5$ and $w=5$ as it seems to be a good compromise between prediction time and recognition performance. Through experiments, we found out that transfer learning from MT-DAN to Meta-DAN leads to better results than from W-DAN or line-level pre-training on both IAM and BRESSAY datasets, so we keep this setting in the following.
@@ 6.3 Meta-DAN
We compare the Meta-DAN with the MT-DAN and the W-DAN for the same hyperparameters ($m=5$ and $w=5$) in Table 6.3. However, it should be noted that this is not a constraint of the proposition, and that $m$ and $w$ can be chosen completely independently for the Meta-DAN. As one can note for the IAM and BRESSAY datasets, combining both windowed queries and multi-token predictions enable to increase the recognition performance, as well as reducing the prediction time. Compared to DAN, the conclusion is not as clear-cut regarding the character error rate as it is in favor of the DAN for IAM with a CER of 4.23% compared to 4.79%, but in favor of the Meta-DAN for BRESSAY with a CER of 2.38% compared to 2.43%. It shows that the Meta-DAN can be both more efficient and faster than the DAN, depending on the target dataset.
@@
Comparison of proposed approaches using the static prediction policy ($m=k$). Approach $m$ $w$ IAM BRESSAY CER Time CER Time DAN 1 1 4.23 3.39 2.43 24.65 W-DAN 1 5 4.87 0.77 3.02 6.63 MT-DAN 5 1 5.14 0.75 2.45 5.28 Meta-DAN 5 5 4.79 0.43 2.38 4.01
@@
Comparison of Meta-DAN ($w=5$, $m=5$) performance with respect to the decoding strategy. Strategy for IAM BRESSAY concurrent CER Time CER Time static ($k=1$) 4.55 0.79 2.31 7.16 static ($k=5$) 4.79 0.43 2.38 4.01 dynamic ($\tau=0.9$) 4.73 0.45 2.37 7.14
Comparing prediction strategies.
As the Meta-DAN relies on multi-token predictions, we can apply the same dynamic prediction policy as for the MT-DAN. This is evaluated in Table 6.3. It appears that the confidence scores associated with the predictions are not as high as those for the MT-DAN, leading to a failure case for the BRESSAY dataset for which the prediction time is close to that of W-DAN (static case with $k=1$). As previously, the static policy with $k=1$ leads to the best recognition performance but suffers from longer prediction. We consider the static policy with $k=m=5$ as the best alternative for the Meta-DAN and use this configuration in the following.
6.4 Visualization